two blocks of masses 2kg and 3kg are connected by massless spring and are placed on smooth surface. tf impulse is given to 2kg block and block starts moving with velocity 10m/s in +ve x direction, then velocity of 2kg block wrt centre of mass is?

Grade:

1 Answers

Aziz Alam IIT Roorkee
8 years ago
The velocity given to the 2 kg block is 10 m/s along x-axis.

The total momentum of the system is P=2*10=20 kg-m/s

This the momentum associated with the centre of mass.
The velocity of the centre of mass= P/M=20/5=4 m/s

So the centre of mass is moving along x-axis with the speed 4 m/s
Then the velocity of the 2 kg block is 6 m/s along the x-axis wrt to centre of mass.